# Partial Resolution of the Erdős-Straus, Sierpiński, and Generalized Erdős-Straus Conjectures Using New Analytical Formulas

### v1

@@ 11.3 Testing Second Formula for Mordell’s Exceptional Cases : between [5000, 10000]
In the following tables, we provide the numerical values of some $n$, $x_{0}$, $y$, and $z$ that satisfy the conditions of each formula. The value $L$ (for "Left") represents the ratio $a/n$, while $R$ (for "Right") is computed as $R=\frac{1}{x_{0}}+\frac{1}{y}+\frac{1}{z}$.
@@
To avoid overloading the article, I have included in the appendix the Python code for each formula so that interested readers can verify for themselves how these formulas generate solutions, especially the second formula, which captures almost all $n$ in a given interval (for $a=4$, the classical Erdős-Straus conjecture), and potentially all $n$, even for very large values, with a significant variety of solutions for each $n$. This can be seen in Tables 2,3,4,5 below, where we capture (for $a=4$) all $n$ in the range $[2,200]$ that satisfy the conjecture using the second formula. I strongly encourage readers to copy and paste the second code from the appendix, run it in a Python environment, and observe how the second formula generates huge solutions (capturing all $n$ for $a=4$ within an interval up to a very high threshold).You will notice when running this code that sometimes a single $n$ can generate around 40 distinct solutions $(x,y,z)$.
I had fun testing naively, numerically (using the fourth code in the appendix of this article) for:
$n\in[2,100],\quad n\in]100,200],\quad n\in]200,300],~{}\cdots,~{}n\in]1900,2000]$
searching for values of $n$ for which the polynomial :
$t^{2}(4x-n)^{2}-2nxt$
admits a perfect square while also identifying the non-captured $n$ (which I call the recalcitrant cases) and computing the percentage of captured $n$ per interval.
I obtained the following results:
For $n\in[2,2000]$ (excluding the subintervals $[500,600],[1200,1300],[1800,1900]$), no recalcitrant cases were found, achieving a 100% capture rate, meaning this polynomial admits a perfect square for all $n$ in these ranges.
For $n\in[500,600]$, a 99.01% capture rate was recorded, with a single recalcitrant case at $n=577$. I handled this manually using the same reasoning proposed at the beginning of this article: Since
$x\geq\left\lceil\frac{577}{4}\right\rceil+1=145,$
I took $x=145$, leading to the discriminant
$\Delta=t(9t-723002).$
Please refer to the fifth code in the appendix of this article to execute this case. You will find that for $t\in[25000,40000]$, we obtain $t=33466$, $\Delta=4479892624$, and $\sqrt{\Delta}=66932$.
For $n\in[1200,1300]$, a 97.03% capture rate was recorded, with three recalcitrant cases:
$n\in\{1201,1249,1277\}.$
For $n\in[1800,1900]$, a 99.01% capture rate was recorded, with a single recalcitrant case at $n=1801$.
For the cases $n\in\{1249,1277\}$, one can proceed in the same way as for $n=577$.
However, the cases $n\in\{1201,1801\}$ are exactly the ones not proven by Mordell and were also not captured in our previous analysis over the range $[2,2000]$, where 15 cases were analyzed, 13 were captured (86%), and the two non-captured cases were also $1201$ and $1801$.
It seems that these two cases require a more refined analysis, which I leave to the mathematical community, as I did not have enough time to focus on them.
(You may also attempt to handle them similarly to $n=577$ by using the fifth code in the appendix to extend the search range for $t$ and $x$. This is just a suggested approach, but I cannot guarantee its effectiveness, as we have infinitely many choices for $t$ and $x$. Therefore, a more detailed analytical study is probably required for these two cases).
Similarly, I had fun using the same code to naively verify whether the polynomial
@@
always admits a perfect square for any fixed $a\geq 5$ and for all $n\geq 2$.
What I observed is that, initially, there are some recalcitrant cases. However, these recalcitrant cases tend to decrease as $n$ increases.
In summary, naively, we observe approximately 95% of captured cases.
However, to reduce the number of recalcitrant cases, the reader may proceed as in the case of $n=577$ previously discussed for $a=4$.
Based on the preceding observations, we numerically observe that the property appears to hold for a very large number of values of $n$. If this regularity is confirmed, one could then formulate it as a conjecture, which we propose here as an open question for the mathematical community.
Conjecture 1: For all $n\in\mathbb{N}$ with $n\geq 2$, there exist $(x_{0},t_{0})\in\mathbb{N}^{*2}$ such that
$t_{0}^{2}(4x_{0}-n)^{2}-2nx_{0}t_{0}=q^{2},\quad q\in\mathbb{N}.$
Conjecture 2: For all $a\in\mathbb{N}$ with $a\geq 5$, and for all $n\in\mathbb{N}$ with $n\geq 2$, there exists $N_{0}\in\mathbb{N}$ such that for all $n\geq N_{0}$, there exist $(x_{0},t_{0})\in\mathbb{N}^{*2}$ such that
$t_{0}^{2}(ax_{0}-n)^{2}-2nx_{0}t_{0}=q^{2},\quad q\in\mathbb{N}.$
@@
Table 6: Results of Sierpiński Conjecture with second formula testing for $n$ from $2$ to $50$
n x y z L R L = R (verification)
3 1 2 6 1.6667 1.6667 True
4 1 8 8 1.25 1.25 True
6 2 4 12 0.8333 0.8333 True
7 5 2 70 0.7143 0.7143 True
8 2 16 16 0.625 0.625 True
9 2 36 36 0.5556 0.5556 True
11 5 4 220 0.4545 0.4545 True
12 3 24 24 0.4167 0.4167 True
13 3 26 78 0.3846 0.3846 True
14 3 84 84 0.3571 0.3571 True
15 4 20 30 0.3333 0.3333 True
16 4 32 32 0.3125 0.3125 True
17 4 34 68 0.2941 0.2941 True
18 4 72 72 0.2778 0.2778 True
19 4 152 152 0.2632 0.2632 True
20 5 40 40 0.25 0.25 True
21 5 42 70 0.2381 0.2381 True
22 5 44 220 0.2273 0.2273 True
23 5 70 322 0.2174 0.2174 True
24 5 240 240 0.2083 0.2083 True
25 6 60 60 0.2 0.2 True
26 6 52 156 0.1923 0.1923 True
27 6 108 108 0.1852 0.1852 True
28 6 168 168 0.1786 0.1786 True
29 6 348 348 0.1724 0.1724 True
30 7 60 140 0.1667 0.1667 True
31 7 62 434 0.1613 0.1613 True
32 7 140 160 0.1562 0.1562 True
33 7 176 336 0.1515 0.1515 True
34 7 476 476 0.1471 0.1471 True
35 8 70 280 0.1429 0.1429 True
36 8 144 144 0.1389 0.1389 True
37 8 148 296 0.1351 0.1351 True
38 8 304 304 0.1316 0.1316 True
39 8 624 624 0.1282 0.1282 True
40 9 120 180 0.125 0.125 True
41 26 12 6396 0.122 0.122 True
42 9 252 252 0.119 0.119 True
43 9 258 774 0.1163 0.1163 True
44 9 792 792 0.1136 0.1136 True
45 10 180 180 0.1111 0.1111 True
46 10 140 644 0.1087 0.1087 True
47 10 188 940 0.1064 0.1064 True
48 10 480 480 0.1042 0.1042 True
49 10 980 980 0.102 0.102 True
50 11 220 220 0.1 0.1 True
@@
Table 7: Results of Sierpiński Conjecture with second formula testing for $n$ from $51$ to $81$
n x y z L R L = R (verification)
51 12 102 204 0.098 0.098 True
52 11 352 416 0.0962 0.0962 True
53 12 106 636 0.0943 0.0943 True
54 11 1188 1188 0.0926 0.0926 True
55 12 220 330 0.0909 0.0909 True
56 12 336 336 0.0893 0.0893 True
57 12 456 456 0.0877 0.0877 True
58 12 696 696 0.0862 0.0862 True
59 12 1416 1416 0.0847 0.0847 True
60 13 260 390 0.0833 0.0833 True
61 95 14 81130 0.082 0.082 True
62 14 124 868 0.0806 0.0806 True
63 13 728 936 0.0794 0.0794 True
64 13 1664 1664 0.0781 0.0781 True
65 14 280 520 0.0769 0.0769 True
66 14 352 672 0.0758 0.0758 True
67 313 14 293594 0.0746 0.0746 True
68 14 952 952 0.0735 0.0735 True
69 14 1932 1932 0.0725 0.0725 True
70 15 420 420 0.0714 0.0714 True
71 15 426 710 0.0704 0.0704 True
72 15 720 720 0.0694 0.0694 True
73 15 876 1460 0.0685 0.0685 True
74 15 2220 2220 0.0676 0.0676 True
75 16 480 480 0.0667 0.0667 True
76 16 608 608 0.0658 0.0658 True
77 16 770 880 0.0649 0.0649 True
78 16 1248 1248 0.0641 0.0641 True
79 16 2528 2528 0.0633 0.0633 True
80 17 400 850 0.0625 0.0625 True
81 18 324 324 0.0617 0.0617 True
@@
Table 8: Results of the generalized conjecture ($a=8$ )with second formula Testing for $n$ from $2$ to $50$
n x y z L R L = R (verification)
5 1 2 10 1.6 1.6 True
6 1 6 6 1.3333 1.3333 True
7 1 14 14 1.1429 1.1429 True
10 2 4 20 0.8 0.8 True
12 2 12 12 0.6667 0.6667 True
13 9 2 234 0.6154 0.6154 True
14 2 28 28 0.5714 0.5714 True
15 2 60 60 0.5333 0.5333 True
16 3 8 24 0.5 0.5 True
18 3 18 18 0.4444 0.4444 True
19 3 12 228 0.4211 0.4211 True
20 3 24 40 0.4 0.4 True
21 3 42 42 0.381 0.381 True
22 3 66 66 0.3636 0.3636 True
23 3 138 138 0.3478 0.3478 True
24 4 24 24 0.3333 0.3333 True
25 4 20 50 0.32 0.32 True
26 5 10 130 0.3077 0.3077 True
27 4 36 54 0.2963 0.2963 True
28 4 56 56 0.2857 0.2857 True
29 4 58 116 0.2759 0.2759 True
30 4 120 120 0.2667 0.2667 True
31 4 248 248 0.2581 0.2581 True
32 5 40 40 0.25 0.25 True
33 5 30 110 0.2424 0.2424 True
34 5 34 170 0.2353 0.2353 True
35 5 70 70 0.2286 0.2286 True
36 5 72 120 0.2222 0.2222 True
37 5 74 370 0.2162 0.2162 True
38 5 190 190 0.2105 0.2105 True
39 5 390 390 0.2051 0.2051 True
40 6 48 80 0.2 0.2 True
41 6 36 1476 0.1951 0.1951 True
42 6 84 84 0.1905 0.1905 True
43 52 6 6708 0.186 0.186 True
44 6 132 132 0.1818 0.1818 True
45 6 180 180 0.1778 0.1778 True
46 6 276 276 0.1739 0.1739 True
47 6 564 564 0.1702 0.1702 True
48 7 56 168 0.1667 0.1667 True
49 7 98 98 0.1633 0.1633 True
50 7 100 140 0.16 0.16 True
Table 9: Results of the generalized conjecture ($a=8$ )with second formula Testing for $n$ from $51$ to $72$
n x y z L R L = R (verification)
51 7 102 238 0.1569 0.1569 True
52 7 140 260 0.1538 0.1538 True
53 8 40 1060 0.1509 0.1509 True
54 7 378 378 0.1481 0.1481 True
55 7 770 770 0.1455 0.1455 True
56 8 112 112 0.1429 0.1429 True
57 8 114 152 0.1404 0.1404 True
58 8 116 232 0.1379 0.1379 True
59 8 118 472 0.1356 0.1356 True
60 8 240 240 0.1333 0.1333 True
61 8 244 488 0.1311 0.1311 True
62 8 496 496 0.129 0.129 True
63 8 1008 1008 0.127 0.127 True
64 9 144 144 0.125 0.125 True
65 9 156 180 0.1231 0.1231 True
66 9 198 198 0.1212 0.1212 True
67 9 134 1206 0.1194 0.1194 True
68 9 204 612 0.1176 0.1176 True
69 9 414 414 0.1159 0.1159 True
70 9 630 630 0.1143 0.1143 True
71 9 1278 1278 0.1127 0.1127 True
72 10 144 240 0.1111 0.1111 True
@@
n x y z L R L = R (verification)
501 46 11523 7682 0.0219560878 0.0219560878 True
503 46 23138 11569 0.0218687873 0.0218687873 True
504 46 46368 15456 0.0218253968 0.0218253968 True
510 49 1785 1666 0.0215686275 0.0215686275 True
513 50 1425 1350 0.0214424951 0.0214424951 True
514 47 24158 12079 0.0214007782 0.0214007782 True
521 48 8336 6252 0.0211132438 0.0211132438 True
523 48 12552 8368 0.0210325048 0.0210325048 True
525 48 25200 12600 0.0209523809 0.0209523809 True
526 48 50496 16832 0.0209125475 0.0209125475 True
528 51 1683 1584 0.0208333333 0.0208333333 True
531 49 7434 5782 0.0207156309 0.0207156309 True
534 49 13083 8722 0.0205992509 0.0205992509 True
536 49 26264 13132 0.0205223881 0.0205223881 True
537 49 52626 17542 0.0204841713 0.0204841713 True
545 51 3706 3270 0.0201834862 0.0201834862 True
546 50 18200 10920 0.0201465201 0.0201465201 True
547 50 27350 13675 0.0201096892 0.0201096892 True
550 51 5610 4675 0.0200000000 0.0200000000 True
556 51 14178 9452 0.0197841727 0.0197841727 True
558 51 28458 14229 0.0197132616 0.0197132616 True
559 51 57018 19006 0.0196779964 0.0196779964 True
567 52 14742 9828 0.0194003527 0.0194003527 True
569 52 29588 14794 0.0193321617 0.0193321617 True
570 52 59280 19760 0.0192982456 0.0192982456 True
575 54 3450 3105 0.0191304348 0.0191304348 True
576 53 10176 7632 0.0190972222 0.0190972222 True
580 53 30740 15370 0.0189655172 0.0189655172 True
581 54 5229 4482 0.0189328743 0.0189328743 True
584 55 3212 2920 0.0188356164 0.0188356164 True
589 54 15903 10602 0.0186757216 0.0186757216 True
590 54 21240 12744 0.0186440678 0.0186440678 True
591 54 31914 15957 0.0186125212 0.0186125212 True
592 54 63936 21312 0.0185810811 0.0185810811 True
594 55 6534 5445 0.0185185185 0.0185185185 True
596 55 8195 6556 0.0184563758 0.0184563758 True
600 55 16500 11000 0.0183333333 0.0183333333 True
601 56 4808 4207 0.0183028286 0.0183028286 True
602 55 33110 16555 0.0182724252 0.0182724252 True
603 55 66330 22110 0.0182421227 0.0182421227 True
609 56 11368 8526 0.0180623974 0.0180623974 True
610 56 13664 9760 0.0180327869 0.0180327869 True
612 65 780 765 0.0179738562 0.0179738562 True
613 56 34328 17164 0.0179445351 0.0179445351 True
620 57 11780 8835 0.0177419355 0.0177419355 True
622 57 17727 11818 0.0176848875 0.0176848875 True
624 57 35568 17784 0.0176282051 0.0176282051 True
@@
[…]
@@
3 def generalized_erdos_straus_conjecture(a, n_min, n_max):

### v2

@@ 13 Testing Second Formula for Mordell’s Exceptional Cases between [5000,10000]
In the following tables, we provide the numerical values of some $n,x_{0},y$, and $z$ that satisfy the conditions of each formula.
The value $L$ (for “Left”) represents the ratio $a/n$, while $R$ (for “Right”) is computed as $R=\frac{1}{x_{0}}+\frac{1}{y}+\frac{1}{z}$ (we use the fractional form here because the values of $n$ tested in the tables below are relatively small, and therefore no floating-point precision issues arise).
@@
According to the equivalence established below between the original equation
$\frac{a}{n}=\frac{1}{x}+\frac{1}{y}+\frac{1}{z}$
and the parametric quadratic equation
$V^{2}-2t(ax-n)V+2nxt=0,$
the case $a=4$ (classical Erdős–Straus conjecture) reassures us that this quadratic equation admits integer solutions at least up to $n=10^{17}$, which is the current numerical verification threshold for the Erdős–Straus conjecture.
For $a=5$ (classical Sierpiński case), we also know today that very high numerical verifications have been carried out. I personally performed quick numerical checks for some values of $a\in\{6,7,8\}$ up to reasonable thresholds (which I do not officially report in this article, as these were rapid tests), and the equation consistently appears to admit integer solutions via the study of its discriminant
$\Delta=4t^{2}(ax-n)^{2}-8nxt=4\bigl(t^{2}(ax-n)^{2}-2nxt\bigr),$
which seems always to contain a perfect square, thus providing integer solutions to the quadratic equation, and by equivalence, to the generalized Erdős–Straus equation.
Since no proof is currently known, these observations lead us to state the following conjectures.
Conjecture 1 (strong version, integer $t$). For every integer $n\geq 2$, there exists $(x,t)\in\mathbb{N}^{*2}$ such that
$t^{2}(4x-n)^{2}-2nxt=q^{2},\qquad q\in\mathbb{N}.$
Conjecture 2 (weak version, rational $t$, not necessarily integer). For every integer $n\geq 2$, there exists $(x,t)\in\mathbb{N}^{*}\times\mathbb{Q}_{+}^{*}$ such that
$t^{2}(4x-n)^{2}-2nxt=q^{2},\qquad q\in\mathbb{Q}_{+}.$
Conjecture 3 (weak version, rational $t$, not necessarily integer). For every fixed integer $a\geq 5$, there exists $N_{0}\geq 2$ in $\mathbb{N}^{*}$ (with $N_{0}$ not strictly smaller than $a/3$) such that for all $n\geq N_{0}$, there exists
$(x,t)\in\mathbb{N}^{*}\times\mathbb{Q}_{+}^{*}\quad\text{with}\quad t^{2}(ax-n)^{2}-2nxt=q^{2},\qquad q\in\mathbb{Q}_{+}.$
Conjecture 4 (strong version, integer $t$). For every fixed integer $a\geq 5$, there exists $N_{1}\geq N_{0}$ ($N_{0}$ is the one given in Conjecture 3), with $N_{1}$ not strictly smaller than $a/3$, such that for all $n\geq N_{1}$, there exists
$(x,t)\in\mathbb{N}^{*2}\quad\text{such that}\quad t^{2}(ax-n)^{2}-2nxt=q^{2},\qquad q\in\mathbb{N}.$
Conjecture 1 is equivalent to the classical Erdős–Straus conjecture. Indeed, under Conjecture 1, the integer $q$ satisfies
$q^{2}=t^{2}(4x-n)^{2}-2nxt,$
and the variables $y$ and $z$ are given by
$y=t(4x-n)-q,\qquad z=t(4x-n)+q.$
It is immediate that $z\in\mathbb{N}^{*}$, since 4xn>04x-n>0 because
$x\geq\left\lfloor\frac{n}{4}\right\rfloor+1>\frac{n}{4},$
and with $t\in\mathbb{N}^{*}$, $x\in\mathbb{N}^{*}$ , $n\in\mathbb{N}^{*}$ and $q\in\mathbb{N}$, we have $z\in\mathbb{N}^{*}$.
Let us now show that $y\in\mathbb{N}^{*}$. From
$q^{2}=t^{2}(4x-n)^{2}-2nxt<t^{2}(4x-n)^{2},$
since 2nxt<0-2nxt<0 for $n,x,t>0$, and since the square root function is increasing on $\mathbb{R}_{+}^{*}$, we obtain
q<t(4xn).q<t(4x-n).
Hence t(4xn)q>0t(4x-n)-q>0, and since $t,x,n,q$ are integers, it follows that
$y=t(4x-n)-q\in\mathbb{N}^{*}.$
We have therefore rigorously shown that whenever the condition
$t^{2}(4x-n)^{2}-2nxt=q^{2},\qquad q\in\mathbb{N},$
is satisfied (Conjecture 1), one can automatically construct $y,z\in\mathbb{N}^{*}$ solving the quadratic equation
$V^{2}-2t(4x-n)V+2nxt=0,$
which, by the equivalence theorem proved earlier, is equivalent to the original Erdős–Straus equation
$\frac{4}{n}=\frac{1}{x}+\frac{1}{y}+\frac{1}{z}.$
Thus, Conjecture 1 is rigorously equivalent to the classical Erdős–Straus conjecture.
Using the same reasoning, one also proves that Conjecture 4 is equivalent to the generalized Sierpiński conjecture ($a\geq 5$) for all $n\geq N_{1}$, where $N_{1}$ is the minimal threshold beyond which the expression
@@
always admits an integer perfect square (under Conjecture 4).
Conjecture 2 is stated for researchers interested in working over $\mathbb{Q}_{+}^{*}$, either to enlarge the search space for $t$, or because it may be useful in other research contexts beyond Egyptian fractions.
Conjecture 3 is particularly useful for numerical verification of the generalized Sierpiński conjecture ($a\geq 5$) for values of $n$ in the finite interval $N_{0},N_{1})$, which is often not too large according to our simulations, since Conjecture 4 addresses the problem for all $n\geq N_{1}$.
Finally, why do we exclude values $n<a/3$? Since $x,y,z\in\mathbb{N}^{*}$, we have
$\frac{1}{x}+\frac{1}{y}+\frac{1}{z}\leq 3,$
which implies $a/n\leq 3$, and hence $n\geq a/3$. Therefore, the decomposition of $a/n$ as a sum of three Egyptian fractions is impossible for all integers $n<a/3$. Consequently, Conjecture 4 cannot hold for such values of $n$, as it is equivalent to the generalized Sierpiński conjecture.
Similarly, even if Conjecture 3 produces rational perfect squares for $n<a/3$, these cannot yield integer solutions to the quadratic equation
$V^{2}-2t(ax-n)V+2nxt=0,$
since, by the equivalence theorem, this quadratic equation admits no integer solutions when $n<a/3$, because the equivalent equation
$\frac{a}{n}=\frac{1}{x}+\frac{1}{y}+\frac{1}{z}$
has no integer solutions in that range.
@@
Table 6: Results of Sierpiński Conjecture with second formula testing for $n$ from 22 to $45$
$n$ $x$ $t$ $y$ $z$ $F$ $q$
2 1 $\tfrac{1}{2}$ 1 2 $\tfrac{1}{4}$ $\tfrac{1}{2}$
3 1 22 2 6 44 22
4 1 88 8 8 0 0
5 2 $\tfrac{4}{5}$ 4 4 0 0
6 2 22 4 12 $16$ 44
7 5 22 2 70 $1156$ $34$
8 2 88 16 16 0 0
9 2 $36$ 36 36 0 0
10 3 $\tfrac{5}{2}$ 10 15 $\tfrac{25}{4}$ $\tfrac{5}{2}$
11 5 88 4 220 $11664$ $108$
12 3 88 24 24 0 0
13 3 $26$ 26 78 $676$ $26$
14 3 $84$ 84 84 0 0
15 4 55 20 30 $25$ 55
16 4 88 32 32 0 0
17 4 $17$ 34 68 $289$ $17$
18 4 $36$ 72 72 0 0
19 4 $152$ 152 152 0 0
20 5 88 40 40 0 0
21 5 $14$ 42 70 $196$ $14$
22 5 $44$ 44 220 $7744$ $88$
23 5 $98$ 70 322 $15876$ $126$
24 5 $240$ 240 240 0 0
25 6 $12$ 60 60 0 0
26 6 $26$ 52 156 $2704$ $52$
27 6 $36$ 108 108 0 0
28 6 $84$ 168 168 0 0
29 6 $348$ 348 348 0 0
30 7 2020 60 140 $1600$ $40$
31 7 $62$ 62 434 $34596$ $186$
32 7 $50$ 140 160 $100$ $10$
33 7 $128$ 176 336 $6400$ $80$
34 7 $476$ 476 476 0 0
35 8 $35$ 70 280 $11025$ $105$
36 8 $36$ 144 144 0 0
37 8 $74$ 148 296 $5476$ $74$
38 8 $152$ 304 304 0 0
39 8 $624$ 624 624 0 0
40 9 $30$ 120 180 $900$ $30$
41 12 $169$ 26 6396 $10144225$ $3185$
42 9 $84$ 252 252 0 0
43 9 $258$ 258 774 $66564$ $258$
44 9 $792$ 792 792 0 0
45 10 $36$ 180 180 0 0
@@
Table 7: Results of Sierpiński Conjecture with second formula testing for $n$ from $46$ to $89$
$n$ $x$ $t$ $y$ $z$ $F$ $q$
46 10 $98$ 140 644 $63504$ $252$
47 10 $188$ 188 940 $141376$ $376$
48 10 $240$ 480 480 0 0
49 10 $980$ 980 980 0 0
50 11 $44$ 220 220 0 0
51 12 $17$ 102 204 $2601$ $51$
52 11 $128$ 352 416 $1024$ $32$
53 11 $512$ 352 1696 $451584$ $672$
54 11 $1188$ 1188 1188 0 0
55 12 $55$ 220 330 $3025$ $55$
56 12 $84$ 336 336 0 0
57 12 $152$ 456 456 0 0
58 12 $348$ 696 696 0 0
59 12 $1416$ 1416 1416 0 0
60 13 $65$ 260 390 $4225$ $65$
61 15 $242$ 66 6710 $11035684$ $3322$
62 14 $62$ 124 868 $138384$ $372$
63 13 $416$ 728 936 $10816$ $104$
64 13 $1664$ 1664 1664 0 0
65 14 $80$ 280 520 $14400$ $120$
66 14 $128$ 352 672 $25600$ $160$
67 15 $134$ 134 2010 $879844$ $938$
68 14 $476$ 952 952 0 0
69 14 $1932$ 1932 1932 0 0
70 15 $84$ 420 420 0 0
71 15 $142$ 426 710 $20164$ $142$
72 15 $240$ 720 720 0 0
73 15 $584$ 876 1460 $85264$ $292$
74 15 $2220$ 2220 2220 0 0
75 16 $96$ 480 480 0 0
76 16 $152$ 608 608 0 0
77 16 $275$ 770 880 $3025$ $55$
78 16 $624$ 1248 1248 0 0
79 16 $2528$ 2528 2528 0 0
80 17 $125$ 400 850 $50625$ $225$
81 18 $36$ 324 324 0 0
82 24 $169$ 52 12792 $40576900$ $6370$
83 20 $125$ 100 4150 $4100625$ $2025$
84 17 $2856$ 2856 2856 0 0
85 18 $240$ 360 2040 $705600$ $840$
86 18 $258$ 516 1548 $266256$ $516$
87 18 $348$ 1044 1044 0 0
88 18 $792$ 1584 1584 0 0
89 18 $3204$ 3204 3204 0 0
@@
Table 8: Results of the generalized conjecture ($a=6$) with second formula, testing for $n$ from 22 to $45$
$n$ $x$ $t$ $y$ $z$ $F$ $q$
2 1 $1/4$ 1 1 0 0
3 1 $2/3$ 2 2 0 0
4 1 2 4 4 0 0
5 1 10 10 10 0 0
6 2 $2/3$ 4 4 0 0
7 3 2 2 42 400 20
8 2 2 8 8 0 0
9 2 4 12 12 0 0
10 2 10 20 20 0 0
11 2 44 44 44 0 0
12 3 2 12 12 0 0
13 3 32 8 312 23104 152
14 3 7 14 42 196 14
15 3 10 30 30 0 0
16 3 24 48 48 0 0
17 3 102 102 102 0 0
18 4 4 24 24 0 0
19 4 32 16 304 20736 144
20 4 10 40 40 0 0
21 4 21 42 84 441 21
22 4 44 88 88 0 0
23 4 184 184 184 0 0
24 5 12 24 120 2304 48
25 5 10 50 50 0 0
26 6 32 16 624 92416 304
27 5 30 90 90 0 0
28 5 70 140 140 0 0
29 5 290 290 290 0 0
30 6 10 60 60 0 0
31 39 2 6 806 160000 400
32 6 24 96 96 0 0
33 6 44 132 132 0 0
34 6 102 204 204 0 0
35 6 420 420 420 0 0
36 7 14 84 84 0 0
37 27 32 8 7992 15936064 3992
38 7 76 76 532 51984 228
39 8 54 36 936 202500 450
40 7 140 280 280 0 0
41 7 574 574 574 0 0
42 8 21 84 168 1764 42
43 8 43 86 344 16641 129
44 8 44 176 176 0 0
45 8 80 240 240 0 0
@@
n x y z L R L = R
501 46 11523 7682 0.021956088 0.021956088 True
503 46 23138 11569 0.021868787 0.021868787 True
504 46 46368 15456 0.021825397 0.021825397 True
510 49 1785 1666 0.021568627 0.021568627 True
513 50 1425 1350 0.021442495 0.021442495 True
514 47 24158 12079 0.021400778 0.021400778 True
521 48 8336 6252 0.021113244 0.021113244 True
523 48 12552 8368 0.021032505 0.021032505 True
525 48 25200 12600 0.020952381 0.020952381 True
526 48 50496 16832 0.020912548 0.020912548 True
528 51 1683 1584 0.020833333 0.020833333 True
531 49 7434 5782 0.020715631 0.020715631 True
534 49 13083 8722 0.020599251 0.020599251 True
536 49 26264 13132 0.020522388 0.020522388 True
537 49 52626 17542 0.020484171 0.020484171 True
545 51 3706 3270 0.020183486 0.020183486 True
546 50 18200 10920 0.020146520 0.020146520 True
547 50 27350 13675 0.020109689 0.020109689 True
550 51 5610 4675 0.020000000 0.020000000 True
556 51 14178 9452 0.019784173 0.019784173 True
558 51 28458 14229 0.019713262 0.019713262 True
559 51 57018 19006 0.019677996 0.019677996 True
567 52 14742 9828 0.019400353 0.019400353 True
569 52 29588 14794 0.019332162 0.019332162 True
570 52 59280 19760 0.019298246 0.019298246 True
575 54 3450 3105 0.019130435 0.019130435 True
576 53 10176 7632 0.019097222 0.019097222 True
580 53 30740 15370 0.018965517 0.018965517 True
581 54 5229 4482 0.018932874 0.018932874 True
584 55 3212 2920 0.018835616 0.018835616 True
589 54 15903 10602 0.018675722 0.018675722 True
590 54 21240 12744 0.018644068 0.018644068 True
591 54 31914 15957 0.018612521 0.018612521 True
592 54 63936 21312 0.018581081 0.018581081 True
594 55 6534 5445 0.018518519 0.018518519 True
596 55 8195 6556 0.018456376 0.018456376 True
600 55 16500 11000 0.018333333 0.018333333 True
601 56 4808 4207 0.018302829 0.018302829 True
602 55 33110 16555 0.018272425 0.018272425 True
603 55 66330 22110 0.018242123 0.018242123 True
609 56 11368 8526 0.018062397 0.018062397 True
610 56 13664 9760 0.018032787 0.018032787 True
612 65 780 765 0.017973856 0.017973856 True
613 56 34328 17164 0.017944535 0.017944535 True
620 57 11780 8835 0.017741935 0.017741935 True
622 57 17727 11818 0.017684887 0.017684887 True
624 57 35568 17784 0.017628205 0.017628205 True
@@
1# Python implementation of the first formula for the Generalized Erdos-Straus Conjecture
@@
3
4def generalized_erdos_straus_conjecture(a, n_min, n_max):